OpenStudy (anonymous):

Differential Equation: If one fundamental solution of (t^2)y''-(3t)y'+3y=0 is y_1(t) = t, what is another fundamental solution y_2(t) satisfying y_2(1)=1 and y'_2(1)=3. (Using the definition of the Wronskian)

OpenStudy (anonymous):

\[t^2y'' - 3ty' + 3y = 0 \] \[y _{1}(t) = t, y _{2}(1)=1, y'_{2}(1)=3. \] Find\[y _{2}(t)\]

OpenStudy (amistre64):

do you know what a wronskian is?

OpenStudy (anonymous):

wronksian, w(y_1,y_2) = y1y2' - y2y1'

OpenStudy (anonymous):

I was thinking that y = C1t + C2e^(RT)

OpenStudy (amistre64):

or are you at the guessing trial and error stage?

OpenStudy (anonymous):

I'm trying to figure out the approach. So yes, im at the trial and error

OpenStudy (anonymous):

i tried to do t^2R^2 -3tR +3 = 0, and find R in terms of T.

OpenStudy (anonymous):

using -b+- sqrt(b^2-4ac) / 2a gives me sqrt(-3t^2) so that did not work as planned.

OpenStudy (amistre64):

let y2 = At, were A is a function of t comes to mind ... but i cant really recall the trial and error stuff

OpenStudy (anonymous):

what if y2 is a complex function.

OpenStudy (amistre64):

y2 = At y2' = A + A't , let A't = 0 y2'' = A' want sure if you were working with complexes

OpenStudy (amistre64):

variation of parameters is what im thinking of

OpenStudy (anonymous):

I was working it in terms of ty2' - y2 and i know this have to be equal of a specific function..

OpenStudy (anonymous):

or result which is not equal to zero

OpenStudy (amistre64):

t^2 A' - 3At -3A't + 3At ------------ 0 = t^2 A' if A' = 0, so let A be a constant maybe?

OpenStudy (amistre64):

y2 = k y1 = t y = y1+y2 y = ct + k

OpenStudy (amistre64):

any thoughts?

OpenStudy (anonymous):

yes...the last method i thought of.

OpenStudy (amistre64):

i cant say im familiar with you last method so im not able to comment on it

OpenStudy (anonymous):

so y1 = t and y2 = e^(RT) ?

OpenStudy (anonymous):

I'm looking up second order homogeneous diff eQ. Actually, we did not cover variation of parameters.

OpenStudy (amistre64):

t^2 r^2 e^(rt) - 3t r e^(rt) + 3 e^(rt) ------------- 0 = e^(rt) (t^2 r^2 -3tr +3)

OpenStudy (anonymous):

this is my approach here....along with Wronskian... http://tutorial.math.lamar.edu/Classes/DE/RealRoots.aspx

OpenStudy (amistre64):

tr = (3 +- sqrt(-3))/2 r = (3 +- sqrt(-3))/2t

OpenStudy (anonymous):

I found something.

OpenStudy (amistre64):

my approach fails the conditions for y2(1) and y2'(1)

OpenStudy (anonymous):

I think i got it.

OpenStudy (amistre64):

does your solution work back inthe setup?

OpenStudy (anonymous):

I used Abel's theorem and based on the equation, the W(y1,y2) = C3/t&+^3 , C3 was used a variable since integrating y2 again will bring forth another variable which is C2 and since I was given y2 and y2' initial condition i used it to substitute and find c2 and c3. I have to try and work backwards.

OpenStudy (anonymous):

it does not :( the w(y1,y2) = 6/t^6 - 4/t it should have been 2/t^3

OpenStudy (amistre64):

it was a heck of a show tho :)

OpenStudy (amistre64):

y2 = t^3 .... try to work towards that

OpenStudy (amistre64):

3t^3 -3t(3t^2) t^2 6t 3t^3 -9t^3 +6t^3 = 0

OpenStudy (anonymous):

I attached the problem just as it is. By the way, how did you deduce y2 to be t^3?

OpenStudy (amistre64):

the wolf ... lol

OpenStudy (amistre64):

what is our def of the wronskian?

OpenStudy (anonymous):

 Must not equal zero for it to have fundamental solution or be linearly independent.